# Enabling Automatic Differentiation with Mollified Graph Neural Operators

### v1

@@ 4.2 Nonlinear Poisson Equation
We consider a nonlinear Poisson equation with varying source terms, boundary conditions, and geometric domain, whose description is detailed in Section D.2. Examples of solutions are shown in Figure 3. The mesh coordinates, signed distance functions, source terms, and boundary conditions, are passed through a GINO model of the form
to approximate the solution $u$. We minimize a weighted sum $\mathcal{L}_{\mathrm{Poisson}}$ (21) of the PDE residual and boundary condition loss. In addition to the results of this section, an ablation study on the choice of GNO radius $r$ and weight function $w$ is provided in Appendix I.
4.2.1 Comparison to Baselines
The results in Figure 4 show the trade-off between inference time and accuracy. $m$GINO achieves a relative squared error 2-3 orders of magnitude lower than Meta-PDE for a comparable running time, and a speedup of 20-25× compared to the solver for similar relative accuracy. In addition, $m$GINO is more consistent across different instances, with smaller variations in errors compared to Meta-PDE.
Table 4: Data loss (MSE) and PDE loss of $m$GINO models trained with different losses $\mathcal{L}$ for Poisson and hyperelasticity equations.
Poisson
Data $1.36\cdot 10^{-5}$ $2.29\cdot 10^{-5}$ $1.39\cdot 10^{-5}$
PDE $3.55\cdot 10^{2}$ $2.07\cdot 10^{-2}$ $7.21\cdot 10^{-3}$
 Hyperelasticity
Data $3.35\cdot 10^{-7}$ $9.69\cdot 10^{-7}$ $9.42\cdot 10^{-5}$
PDE $2.97\cdot 10^{26}$ $1.57\cdot 10^{-2}$ $1.21\cdot 10^{-2}$
(a)(b)(c)
Figure 5: (a) Prediction of a $m$GINO model trained using data loss only for the nonlinear Poisson equation. This is the predicted solution evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. We see that the prediction exhibits discontinuities that coincide with the circles of radius $r$ (blue lines) centered at the neighboring latent query points. (b)(c) Norm of the gradient of the predicted solution shown in (a), computed using automatic differentiation (in (b)) and finite differences (in (c)).
4.2.2 Using Different Training Losses
We trained models using a data loss, a physics loss, and a hybrid loss (data and physics). Table 4 shows that using a physics loss has a comparable data error to the data-driven approach, while achieving a PDE residual 4-5 orders of magnitude lower. The hybrid approach proved more challenging to tune and did not perform as well.
When training only with data loss, the predicted solutions have discontinuities at higher resolutions coinciding with circles of radius $r$ centered at the latent query points of the GINO, (see Figure 5(a)). These discontinuities lead to high derivatives and inaccurate physics losses, regardless of the differentiation method used (see Figures 5(b)(c)). This happens despite the low data MSE, indicating that only training the $m$GINO model with data loss is not sufficient to capture the solution correctly at higher resolutions.
Recall that the GNO’s kernel integration can be viewed as an aggregation of messages if we construct a graph on the spatial domain of the PDE, as described in Li et al. (2020b).
The mean aggregation is given by
where $v_{t}(x)\in\mathbb{R}^{n}$ are the node features, $e(x,y)\in\mathbb{R}^{n_{e}}$ are the edge features, $W\in\mathbb{R}^{n\times n}$ is learnable, $N(x)$ is the neighborhood of $x$, and $\kappa_{\phi}(e(x,y))$ is a neural network mapping edge features to a matrix in $R^{n\times n}$. When using a differentiable weighting function, points in $N(x)$ at the edge of the neighborhood have near-zero weights but still contribute to the denominator $|N(x)|$. Thus, as the query point $x$ moves slightly, additional neighbors get included or excluded with near-zero weights, thereby introducing the discontinuities we see in Figure 5(a). Using a sum aggregation for the output GNO’s kernel integration mitigates the rigid patterns. When training with a physics loss, the patterns disappear, as shown in Figure 16, while the MSE for the predictions remains of the same order of magnitude.
Remark 4.1.
Finetuning the trained data-driven model using $\mathcal{L}_{\text{Poisson}}$ did not work well. The data-driven model has a physics loss 7 orders of magnitude larger than the data loss, and attempts at lowering the physics loss only worked by completely sacrificing the data loss.
4.2.3 Poisson Loss with Finite Differences (FD)
(a)(b)(c)
(d)(e)(f)
Figure 6: Prediction of a $m$GINO model and the norm of its gradient, computed on a $16\times 16$ regular grid (top row) and a $96\times 96$ regular grid (bottom row) using autograd and finite differences (FD). All plots are displayed on a $16\times 16$ grid for ease of qualitative comparison.
Figure 6 shows gradients norms on a domain patch, computed using autograd and FD at $16\times 16$ and higher resolutions (both displayed at $16\times 16$). The FD gradient norms at lower resolution in (c) differ completely from those obtained using autograd (b)(e) and high-resolution FD (f), while the latter are very similar. In addition, inaccuracies can be further amplified with FD on unstructured point clouds in regions with a low density of points. The autograd $m$GINOs were trained successfully with a density of randomly located points on the whole domain slightly lower than the density here with $16^{2}$ points on the domain patch displayed in Figure 6. Given that numerical errors made on derivatives are amplified in physics losses, $\mathcal{L}_{\text{Poisson}}$ does not provide enough information to move towards physically plausible solutions when computed using FD on unstructured point clouds at the same resolution at which autograd $m$GINOs models were trained successfully. This showcases how autograd $m$GINO can be used to obtain surrogate models using physics losses at resolutions at which FD is not accurate enough. We estimate that at least 9× more points would be necessary to compute reasonable FD derivatives. Higher resolution FD computations typically require a higher computational time and memory requirement (see Appendix H for a comparison of training times at different resolutions).
4.3 Hyperelasticity Equation
We consider a hyperelasticity equation modeling the deformation of a 2D porous hyperelastic material under compression with varying pore sizes, whose description is detailed in Section D.3. The mesh coordinates and signed distance functions are passed through a GINO model of the form
to approximate the solution. This solution can be obtained as the minimizer of the total Helmholtz free energy of the system, which we use as our loss function instead of the PDE loss from the strong form of the hyperelasticity equation. We also add a weighted boundary loss term.
We trained models using data and/or physics losses. Table 4 shows that using a physics loss is critical for this problem. The data-driven approach achieves a low data loss, but the predictions do not satisfy the physics at all. In contrast, both the physics only and hybrid approaches achieve low data and PDE errors, with the hybrid approach performing the best. As for the Poisson equation, finetuning the trained data-driven model using the physics loss did not work, due to the high physics loss of the data-driven model. As in Section 4.2.3, autograd allowed to compute accurate derivatives at resolutions where finite differences are not accurate.
A comparison to baselines is displayed in Figure 4. $m$GINO achieves a relative squared error 2 orders of magnitude lower than Meta-PDE with a slightly faster running time, and achieves a speedup of 3000-4000× compared to the FEniCS solver for similar accuracy. In addition, $m$GINO achieves consistent results across different samples (all within a single order of magnitude) while Meta-PDE results span more than 3 orders of magnitude. This high variation in Meta-PDE results is likely caused by the difficulty disparity across samples, as samples with larger pores are harder to resolve.
Figure 7: The airfoil design problem. Parametrized vertical displacements of control spline nodes generate a point cloud, which is passed through the differentiable $m$GINO to obtain a pressure field $p$, from which we can compute the drag-lift ratio. We update the control nodes to minimize the drag-lift ratio by differentiating through this entire procedure.
4.4 Airfoil Inverse Design
We consider the transonic flow over an airfoil, governed by the Euler equation. An initial NACA-0012 shape is mapped onto a ‘cubic’ design element with $8$ control nodes in the vertical direction. That initial shape is morphed to a different shape following the displacement field of the control nodes. A more detailed description is provided in Section D.3
For the forward problem, the mesh point locations and signed distance functions are passed through a differentiable mollified GINO model of the form
to produce an approximation of the pressure field $p$, achieving a validation (relative L2) data loss of $0.0146$ (trained in a data-driven manner). We use this trained model for inverse design. More precisely, we parametrize the shape of the airfoil by the vertical displacements of a few spline nodes, and set the design goal to minimize the drag-lift ratio. The parametrized displacements of the spline nodes are mapped to a mesh, which is passed through the $m$GINO to obtain a pressure field, from which we can obtain the drag lift ratio. We optimize the vertical displacement of spline nodes by differentiating through this entire procedure. A depiction of the airfoil design problem is given in Figure 7.
As a result of this optimization process, we obtain an airfoil design with drag coefficient 0.0216 and lift coefficient 0.2371, based on the model prediction. Using the numerical solver on this optimal design, we verify these predictions and obtain a similar drag coefficient 0.0217 and a similar lift coefficient 0.2532. This yields a drag-lift ratio of around 0.09, outperforming the optimal drag-lift ratio of 0.14 reported by Li et al. (2022) (drag 0.04 and lift 0.29, obtained using a Geo-FNO instead of $m$GINO).
Discussion
We proposed $m$GNO, a fully differentiable version of GNO, to allow for the use of automatic differentiation when computing derivatives, and embedded it within GINO to learn efficiently solution operators of families of large-scale PDEs with varying geometries without data. The proposed approach circumvents the computational limitations of traditional solvers, the heavy data requirement of fully data-driven approaches, and the generalization issues of PINNs.
The use of a physics loss proved critical in most experiments, and was sufficient to achieve good results in the absence of data. This highlights the need for efficient and accurate methods to compute physics losses, to improve data-efficiency and regularize neural operators. Autograd can compute exact derivatives in a single pass seamlessly across complex geometries and enables higher-order derivatives with minimal additional cost. However, it can be memory-intensive for deep models, possibly making it less efficient than finite differences (FDs) for simpler problems. Despite these limitations, its accuracy, capability to handle complex geometries, and scalability for complicated learning tasks make it the preferred differentiation method in PINNs, and a promising approach in our physics-informed neural operator setting on complex domains.
Using autograd instead of FDs led to a 20× reduction of the relative L2 data loss for Burgers’ equation on regular grids, suggesting that the autograd physics loss better captured the physics. underlying the data Autograd $m$GINO performed seamlessly for the Poisson and hyperelasticity equations on unstructured point clouds, while FDs were not sufficiently accurate at the training resolution used and would need at least 9× more points to compute reasonable derivatives. Autograd $m$GINO achieved a relative error 2-3 orders of magnitude lower than the Meta-PDE baselines for a comparable running time, and enjoyed speedups of 20-25× and 3000-4000× compared to the solver for similar accuracy on the Poisson and hyperelasticity equations. We demonstrated with an airfoil design problem that $m$GINO can be used seamlessly for solving inverse design and shape optimization problems on complex geometries.
In the future, we intend to investigate the use of adaptive loss balancing schemes for hybrid losses. Other mechanisms can improve the performance of PINNs, and we plan to explore their integration with the proposed approach. The training computational cost could also be amortized, for instance via higher-order automatic differentiation method such as the Stochastic Taylor Derivative Estimator (Shi et al., 2024).
Code
The PyTorch codes used for experiments presented in this paper have been added to the open-source neural operator library from Kossaifi et al. (2024) at https://github.com/neuraloperator/neuraloperator.
Acknowledgments
Ryan Lin is supported by the Caltech Summer Undergrad- uate Research Fellowships (SURF) program. David Pitt is supported by the Schmidt Scholars in Software Engineering program. Anima Anandkumar is supported in part by Bren endowed chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 senior fellow program at Schmidt Sciences.
Appendix A FNO and GINO architectures
Figure 8: The Fourier Neural Operator (FNO) architecture (extracted from (Li et al., 2020a)).
Figure 9: The Geometry-Informed Neural Operator (GINO) architecture (extracted from (Li et al., 2023)).
Appendix B Weighting Functions
Figure 10: Examples of weighting functions (9)-(12) for $m$GNO.
Appendix C $m$GNO Layer Pseudocode
Here, we provide a simplified example of PyTorch pseudocode for the $m$GNO layer
with the half_cos weighting function
$w_{\text{half\_cos}}(x,y)=\mathds{1}_{\text{B}_{r}(x)}(y)\left[0.5+0.5\cos{\!\left(\pi d\right)}\right],$ (15)
where $d=\|x-y\|^{2}/r^{2}$.
Appendix D Problems Considered
D.1 Burgers’ Equation
$u(x,0)$ $=u_{0}(x).$
We focus on the dataset of Li et al. (2021a, b) consisting of 800 instances of the Burgers’ equation with viscosity coefficient $\nu=0.01$ and $128\times 26$ resolution. Each sample in the dataset corresponds to a different initial condition $u_{0}$ drawn from the Gaussian process $\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Examples of solutions are shown in Figure 11.
$v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ (16)
D.2 Nonlinear Poisson Equation
@@
$\nabla\cdot\left[(1+0.1u(\textbf{x})^{2})\nabla u(\textbf{x})\right]$ $=f(\textbf{x})$ $\textbf{x}\in\Omega$ (18)
$u(\textbf{x})$ $=b(\textbf{x})$ $\textbf{x}\in\partial\Omega$ (19)
where $u\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{2}$.
The domain $\Omega$ is centered at the origin and defined in polar coordinates with varying radius about the origin
@@
where the parameters $c_{1}$ and $c_{2}$ are drawn from a uniform distribution on $(-0.2,0.2)$. The source term $f$ is a sum of radial basis functions $f(\textbf{x})=\sum_{i=1}^{3}\beta_{i}\exp{||\textbf{x}-\mu_{i}||_{2}^{2}},$ where $\beta_{i}\in\mathbb{R}$ and $\mu_{i}\in\mathbb{R}^{2}$ are both drawn from standard normal distributions. The boundary condition $b$ is a periodic function, defined in polar coordinates as $b_{0}+\frac{1}{4}[b_{1}\cos(\theta)+b_{2}\sin(\theta)+b_{3}\cos(2\theta)+b_{4}\sin(2\theta)]$, where the parameters $b_{i}$ are drawn from a uniform distribution on $(-1,1)$. This is the setting used by Qin et al. (2022).
@@
Examples of predictions made by the trained $m$GINO model for a variety of geometries are displayed in Figure 12, below the corresponding reference solutions.
@@
Figure 12: Comparison of solutions to nonlinear Poisson equations (top) with the corresponding $m$GINO predictions (bottom).
D.3 Hyperelasticity Equation

### v2

@@ 4.2 Nonlinear Poisson Equation
4.2.1 Problem Description
@@
$\nabla\cdot\left[(1+0.1u(\textbf{x})^{2})\nabla u(\textbf{x})\right]$ $=f(\textbf{x})\ \$$\textbf{x}\in\Omega$ (16)
$u(\textbf{x})$ $=b(\textbf{x})\$$\textbf{x}\in\partial\Omega$ (17)
where $u\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{2}$. The domain $\Omega$ is centered at the origin and defined in polar coordinates with varying radius about the origin
@@
[…]
@@
Examples of predictions made by the trained mmGINO model for a variety of geometries are displayed in Figure˜6, below the corresponding reference solutions.
@@
Figure 6: Comparison of reference solutions to the nonlinear Poisson equation on various geometries (top) with the corresponding mmGINO predictions (bottom).
4.2.2 Comparison to Baselines
The results in Figure˜7 show the trade-off between inference time and accuracy. mmGINO achieves a relative squared error 2-3 orders of magnitude lower than Meta-PDE (i.e. initialized PINNs) for a comparable running time, and a speedup of 20-25× compared to the solver for similar relative accuracy. In addition, mmGINO is more consistent across different instances, with smaller variations in errors compared to Meta-PDE.
4.2.3 Using Different Training Losses
We trained models using a data loss, a physics loss, and a hybrid loss (data and physics). Table˜6 shows that using a physics loss has a comparable data error to the data-driven approach, while achieving a PDE residual 4-5 orders of magnitude lower. Fine-tuning the trained data-driven model using $\mathcal{L}_{\text{Poisson}}$ did not work well, as its physics loss is 7 orders of magnitude higher than its data loss, and attempts at lowering the physics loss only worked by completely sacrificing the data loss.
Table 6: Data and PDE losses of mmGINO trained with different losses $\mathcal{L}$ for the Poisson equation.
   MSE Data Loss $1.36\cdot 10^{-5}$ $2.29\cdot 10^{-5}$ $1.39\cdot 10^{-5}$
   PDE Residual Loss $3.55\cdot 10^{2}$ $2.07\cdot 10^{-2}$ $7.21\cdot 10^{-3}$
When training only with data loss, the predicted solutions have discontinuities at higher resolutions coinciding with circles of radius $r$ centered at the latent query points of the GINO (see Figure˜8(a)). These discontinuities lead to high derivatives and inaccurate physics losses, regardless of the differentiation method used (see Figures 8(b)(c)). This happens despite the low data MSE, indicating that only training the mmGINO model with data loss is not sufficient to capture the solution correctly at higher resolutions.
Recall that the GNO’s kernel integration can be viewed as an aggregation of messages if we construct a graph on the spatial domain of the PDE, as described in Li et al. (2020b). The mean aggregation is given by
where $v_{t}(x)\in\mathbb{R}^{n}$ are the node features, $e(x,y)\in\mathbb{R}^{n_{e}}$ are the edge features, $W\in\mathbb{R}^{n\times n}$ is learnable, $N(x)$ is the neighborhood of $x$, and $\kappa_{\phi}(e(x,y))$ is a neural network mapping edge features to a matrix in $R^{n\times n}$. When using a differentiable weight function, points in $N(x)$ at the edge of the neighborhood have near-zero weights but still contribute to the denominator $|N(x)|$. Thus, as the query point $x$ moves slightly, additional neighbors get included or excluded with near-zero weights, thereby introducing the discontinuities we see in Figure˜8(a). Using a sum aggregation for the output GNO’s kernel integration mitigates the rigid patterns. When training with a physics loss, the patterns disappear, as shown in Figure˜9, while the MSE for the predictions remains of the same order of magnitude.
(a)(b)(c)
Figure 8: (a) Prediction of a mmGINO trained using data loss only for the nonlinear Poisson equation. This is the predicted solution evaluated at a high resolution on a small patch centered at a latent query point of the mmGNO. The prediction exhibits discontinuities that coincide with the circles of radius $r$ (blue lines) centered at the neighboring latent query points. (b)(c) Norm of the gradient of the predicted solution shown in (a), computed using automatic differentiation (in (b)) and finite differences (in (c)).
Figure 9: Comparison between the prediction of a mmGINO model (left) trained using a hybrid loss, $\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ and the corresponding ground truth solution (right) for the nonlinear Poisson equation. The prediction and ground truth solutions are evaluated at a high resolution on a small patch centered at a latent query point of the mmGNO. Using the physics loss regularized the higher resolution prediction, and removed the visible discontinuity patterns shown in Figure˜8(a).
4.2.4 Nonlinear Poisson PDE Residual Loss with Finite Differences
Figure˜10 shows gradients norms on a domain patch, computed using Autograd and finite differences at $16\times 16$ and higher resolutions (both displayed at $16\times 16$). The finite difference gradient norms at lower resolution in (c) differ completely from those obtained using Autograd (b)(e), and high-resolution finite differences (f), while the latter are very similar. In addition, inaccuracies can be further amplified with finite differences on unstructured point clouds in regions with a low density of points. The Autograd mmGINOs were trained successfully with a density of randomly located points on the whole domain slightly lower than the density here with $16^{2}$ points on the domain patch displayed in Figure˜10. Given that numerical errors made on derivatives are amplified in physics losses, $\mathcal{L}_{\text{Poisson}}$ does not provide enough information to move towards physically plausible solutions when computed using finite differences on unstructured point clouds at the same resolution at which Autograd mmGINOs models were trained successfully. This showcases how Autograd mmGINO can be used to obtain surrogate models using physics losses at resolutions at which finite differences are not accurate enough. We estimate that at least 9× more points would be necessary to compute reasonable finite difference derivatives. Higher resolution finite difference computations typically require a higher computational time and memory requirement (see Appendix˜F for a comparison of training times at different resolutions).
(a)(b)(c)
(d)(e)(f)
Figure 10: Prediction of a mmGINO model and the norm of its gradient, computed on a $16\times 16$ regular grid (top row) and a $96\times 96$ regular grid (bottom row) using Autograd and finite differences (FD). All plots are displayed on a $16\times 16$ grid for ease of qualitative comparison.
4.2.5 Data Efficiency
We now investigate the data efficiency of hybrid loss training, and more precisely how performance changes when varying the number of points at which the data loss and physics loss are evaluated when using a hybrid training loss $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ for the nonlinear Poisson equation. Table˜7 displays the results obtained with the proposed mmGINO model, showing that both the data and physics losses worsen when the number of points used for the physics loss during training decreases. However, performance is not very heavily impacted when the number of points used for the data loss during training is reduced.
Table 7: Data loss (MSE) and physics loss (computed using Autograd) for the proposed mmGINO models on the Poisson equation, when training with a hybrid loss. We vary the resolutions at which the data loss and physics losses are evaluated during training. Each row corresponds to a different number of points used for the physics loss, while each column corresponds to a different number of points used for the data loss.
Physics Data $400$ points $200$ points $50$ points
Physics $|$ Data Physics $|$ Data Physics $|$ Data
$800$ points $2.63\cdot 10^{-2}\ |\ 8.28\cdot 10^{-5}$ $3.31\cdot 10^{-2}\ |\ 3.89\cdot 10^{-5}$ $3.82\cdot 10^{-2}\ |\ 6.73\cdot 10^{-5}$
$400$ points $3\cdot 10^{-1}\ |\ 1.10\cdot 10^{-4} 1.221$$2\cdot 10^{-1}\ |\ 9.20\cdot 10^{-5}$ $1.39\cdot 10^{-1}\ |\ 1.29\cdot 10^{-4}$
$200$ points $3.75\cdot 10^{-1}\ |\ 2.40\cdot 10^{-4}$ $2.55\cdot 10^{-1}\ |\ 2.92\cdot 10^{-4}$ $3.33\cdot 10^{-1}\ |\ 5.18\cdot 10^{-4}$
4.2.6 Ablation Study on GNO Radius and Weight Function
We investigate how performance changes as the mmGNO radius changes for mmGINO models, and also compare the performance of the different weight functions (9)-(12) on the nonlinear Poisson equation. Note that the radius cutoff used in the weight functions of mmGNOs is the same as the mmGNO radius to avoid the additional cost of carrying a second neighbor search. The mmGNO radius $r$ is an important hyperparameter to tune. It needs to be large enough so that the ball $\text{B}_{r}(x)$ contains sufficiently many other latent query points, and a value too large can lead to prohibitive computational and memory costs. We denote the number of latent query points in $\text{B}_{r}(x)$ by $|\text{B}_{r}(x)|$. For the nonlinear Poisson equation, we are using a regular 2D latent space grid, so $\text{B}_{r}(x)$ will only contain a single latent query point ($x$ itself) until $r$ is large enough for $\text{B}_{r}(x)$ to contain one extra latent query point in each direction. $\text{B}_{r}(x)$ will progressively contain more latent query points by thresholds as the radius increases further, together with computational time and memory cost.
Table˜8 shows the results obtained for mmGINO models with different weight functions for values of $r$ such that $|\text{B}_{r}(x)|=3^{2},5^{2},7^{2},9^{2},15^{2}$. As expected, performance is poor when the radius is too small. On the other hand, when $r$ becomes too large, the computational and memory costs significantly increase while performance deteriorates. For this numerical experiment, the best performance was achieved with a radius for which $|\text{B}_{r}(x)|=7^{2}$, and $w_{\text{half\_cos}}$ and $w_{\text{quartic}}$ led to the best performance.
Table 8: MSE of mmGINO models trained on only PDE loss, with different GNO radii and weight functions.
bump half_cos quartic octic
$|\text{B}_{r}(x)|=9$ ($r=0.0875$) $3.64\cdot 10^{-1}$ $3.41\cdot 10^{-1}$ $2.20\cdot 10^{-2}$ $3.48\cdot 10^{-1}$
$|\text{B}_{r}(x)|=25$ ($r=0.13125$) $4.51\cdot 10^{-5}$ $3.59\cdot 10^{-5}$ $3.95\cdot 10^{-5}$ $4.24\cdot 10^{-5}$
$|\text{B}_{r}(x)|=49$ ($r=0.175$) $1.40\cdot 10^{-4}$ $\mathbf{1.39\cdot 10^{-5}}$ $2.93\cdot 10^{-5}$ $9.49\cdot 10^{-5}$
$|\text{B}_{r}(x)|=81$ ($r=0.21875$) $5.35\cdot 10^{-4}$ $1.01\cdot 10^{-4}$ $7.79\cdot 10^{-5}$ $1.76\cdot 10^{-4}$
$|\text{B}_{r}(x)|=225$ ($r=0.35$) $3.92\cdot 10^{-3}$ $9.36\cdot 10^{-4}$ $4.28\cdot 10^{-3}$ $6.59\cdot 10^{-4}$
4.3 Hyperelasticity Equation
4.3.1 Problem Description